# D³-Predictor: Noise-Free Deterministic Diffusion for Dense Prediction

### v3

@@ Abstract
Although diffusion models with strong visual priors have emerged as powerful dense prediction backboens, they overlook a core limitation: the stochastic noise at the core of diffusion sampling is inherently misaligned with dense prediction that requires a deterministic mapping from image to geometry. In this paper, we show that this stochastic noise corrupts fine-grained spatial cues and pushes the model toward timestep-specific noise objectives, consequently destroying meaningful geometric structure mappings. To address this, we introduce D³-Predictor, a noise-free deterministic framework built by reformulating a pretrained diffusion model without stochasticity noise. Instead of relying on noisy inputs to leverage diffusion priors, D³-Predictor views the pretrained diffusion network as an ensemble of timestep-dependent visual experts and self-supervisedly aggregates their heterogeneous priors into a single, clean, and complete geometric prior. Meanwhile, we utilize task-specific supervision to seamlessly adapt this noise-free prior to dense prediction tasks. Extensive experiments on various dense prediction tasks demonstrate that D³-Predictor achieves competitive or state-of-the-art performance in diverse scenarios. In addition, it requires less than half the training data previously used and efficiently performs inference in a single step. Our code, data, and checkpoints are publicly available at https://x-gengroup.github.io/HomePage_D3-Predictor/.
@@ 1 Introduction
Dense prediction, such as depth estimation [ming2021deep] and surface normal estimation [wang2015designing], are fundamental tasks in computer vision, with numerous applications like autonomous driving [cabon2020virtual, godard2019digging], scene reconstruction [chen2024nc, zhang2025monoinstance], inverse rendering [chen2024intrinsicanything, zeng2024rgbx], and so on. Although state-of-the-art discriminative dense prediction models [bochkovskii2024depth, hu2024metric3d] achieve impressive performance, they still struggle to capture fine-grained high-frequency details. To address this limitation, current works [ke2024repurposing, lee2024exploiting, fu2024geowizard, ye2024stablenormal, zhao2025diception] reformulate dense prediction as an image-conditioned iterative denoising process based on diffusion models [rombach2022high]. By leveraging powerful visual priors of diffusion models, these methods can produce dense predictions results with fine-grained geometric details.
@@
While these diffusion-based dense prediction methods have demonstrated promising results, they still suffer from the stochastic noise inherent in diffusion models. Stochastic noise is a fundamental component of diffusion models, enabling sample diversity that is particularly beneficial for creative image [li2024photomaker, xu2024diffusion, qin2023unicontrol] and video [yang2024cogvideox, zhao2024moviedreamer, wang2024framer] generation. However, dense prediction is intrinsically deterministic, indicating that the stochastic noise essential to diffusion models may be misaligned with the task’s deterministic objective. We posit that the stochastic noise introduces two critical issues for dense prediction tasks: 1) Stochastic noise disrupts the geometric structures and small-scale objects in the input image, thereby degrading the integrity of input information and impeding precise spatial perception (cf. Fig. 2 (a)); 2) The stochastic noise drives diffusion models to focus on modeling noise distributions [ho2020denoising] instead of establishing the geometric structure mappings essential for dense prediction tasks [fu2024geowizard]. Moreover, the iterative denoising process of the diffusion model further incurs substantial inference overhead. These observations motivate a fundamental question: Can diffusion models be reformulated into a noise-free deterministic framework to better suit dense prediction tasks?
@@
Recent works attempt to employ deterministic noise to alleviate the stochasticity in diffusion-based dense prediction methods. For example, GenPercept [xu2024matters] and E2E-FT [garcia2025fine] eliminate stochasticity by fixing the noise schedule (a function of the timestep) to introduce deterministic noise. However, diffusion models learn timestep-specific objectives [balaji2022ediff, biroli2024dynamical], leading to distinct diffusion priors at each timestep. Consequently, fixing the noise schedule disrupts this prior structure, resulting in incomplete priors and a loss of geometric fidelity (cf. Fig. 2 (b)). On the other hand, StableNormal [ye2024stablenormal] suppresses stochasticity while preserving diffusion priors via a more complex two-stage pipeline with external DINO [oquab2023dinov2] guidance. However, this design incurs a higher computational overhead.
@@
In this work, we aim to fully eliminate the adverse effects of stochastic noise on diffusion-based dense prediction, without compromising the diffusion prior and with minimal additional computational cost. To this end, we propose D³-Predictor, a noise-free deterministic diffusion-based dense prediction model initialized from the pretrained diffusion model. Specifically, we treat the pretrained diffusion models at different timesteps as an ensemble of visual experts, each exhibiting distinct timestep-dependent diffusion priors. In this context, each visual expert takes a noisy image together with its timestep as input, while our D³-Predictor operates directly on a clean image. D³-Predictor then aggregates diffusion priors from multiple visual experts into a complete and noise-free one in a lightweight self-supervised manner, by aligning its internal representations with those of visual experts. We further apply task-specific supervision to the D³-Predictor to easily leverage this aggregated diffusion prior for various dense prediction tasks.
@@
  • We introduce D³-Predictor, a noise-free deterministic diffusion-based dense prediction model. The key insight of D³-Predictor lies in leveraging self-supervised learning to obtain noise-free and complete diffusion priors, which can be effectively utilized for dense prediction through simple task-specific supervision.
@@ 2.1 Diffusion-based Dense Prediction
Several works have explored leveraging diffusion models for dense prediction. Methods such as DIFT [tang2023emergent] and CleanDIFT [stracke2025cleandift] extract diffusion priors as general-purpose visual features and train lightweight probes for dense prediction tasks, but they are not end-to-end and their performance remains limited. While CleanDIFT [stracke2025cleandift] treats the diffusion model as a series of feature extractors, our method views it as an ensemble of visual experts directly optimized for dense prediction via end-to-end, task-specific supervision. On the other hand, other methods [ke2024repurposing, gui2025depthfm, zhao2025diception, fu2024geowizard] fine-tune a pretrained diffusion model for dense prediction tasks, achieving impressive results. However, these fine-tuning methods rely on stochastic noise to exploit diffusion priors. GenPercept [xu2024matters] and E2E-FT [garcia2025fine] attempt to alleviate stochasticity but disrupt the integrity of diffusion priors, while StableNormal [ye2024stablenormal] suppresses stochasticity at the cost of additional computational overhead. In contrast, our method reformulates the diffusion model into a noise-free framework to better suit dense prediction in a simple self-supervised manner, ensuring both prediction performance and speed.
@@ 2.2 Self-Supervised Enhanced Diffusion Models
Recent studies [yu2024representation, wu2025representation, jiang2025no] have shown that enhancing the internal representations of diffusion models, which serve as diffusion priors, can accelerate generative training and improve generation quality. Unlike methods [yu2024representation] that rely on external pretrained features (e.g., DINO [oquab2023dinov2]), SRA [jiang2025no] improves diffusion representations in a self-supervised way via self-representation alignment. The self-supervised enhancement in SRA leads to improvements in both image generation quality and efficiency, demonstrating that diffusion models can provide representation guidance by themselves. In this work, we extend the idea of this self-supervised enhancement from diffusion-based generation to diffusion-based dense prediction. We align the noise-free internal representations of the diffusion model with their noisy counterparts at different noise levels in a self-supervised manner. This process produces complete, noise-free diffusion priors that are inherently more suitable for dense prediction.
@@ 3.1 Preliminary: Diffusion-based Dense Prediction
As illustrated in Fig. 3 (b), previous diffusion-based dense prediction methods, such as Marigold [ke2024repurposing], adopt the time-consuming iterative denoising process described above, where the conditional input $c$ is an RGB image. At each iteration, the model progressively removes stochastic noise associated with a specific noise level and timestep, thus necessitating multi-step inference.
In contrast, recent work GenPercept [xu2024matters] simplifies this process by fixing the noise schedule. However, it is worth noting that diffusion models have different objectives at different noise levels [balaji2022ediff, biroli2024dynamical]. At higher noise levels, the model captures coarse structural information, while at lower noise levels, it focuses on details [rissanen2022generative]. Thus, the model learns distinct diffusion priors across timesteps, which together constitute a complete diffusion prior. Fixing the noise schedule disrupts this prior structure, leading to incomplete priors and suboptimal performance.
@@ 3.2.1 Self-Supervised Diffusion Priors Aggregation
The core of our approach is to train our D³-Predictor $f_{D^{3}}(\cdot)$ to incorporate a noise-free and complete diffusion prior compared to the pre-trained diffusion model, which is enhanced to be more suitable for deterministic dense prediction tasks. We achieve this enhancement through a simple self-supervised manner, and crucially, relying solely on the model itself rather than costly external guidance (e.g., DINO [oquab2023dinov2]) adapted by previous works [ye2024stablenormal]. We begin by initializing D³-Predictor with a pretrained diffusion model.
As shown in Fig 3 (a), the insight here is to view the diffusion model as an ensemble of $t$ visual experts, where each expert $f_{\text{exp}}(\cdot,\epsilon,t)$ exhibits a distinct timestep-dependent diffusion prior. We inherit the mechanism of aligning a clean backbone with a frozen diffusion model from CleanDIFT [stracke2025cleandift]. While the visual experts take a noisy image $(\epsilon,x)$ together with its timestep $t$ as input, our D³-Predictor operates directly on a clean image $x$. During training, D³-Predictor aligns its internal representations $r_{D^{3}}^{K}$ = $f_{D^{3}}^{K}(x)$ with those of $t$ visual experts $r_{\text{exp}}^{K,t}$ = $f_{\text{exp}}^{K}(x,\epsilon,t)$. Here, $K$ is the index set of model layers selected for alignment.
To stabilize this timestep-dependent representation alignment, we introduce a lightweight, trainable timestep-conditioned projection head $P_{\varphi}^{K}(\cdot,t)$ parameterized by $\varphi$. While the use of projection heads for representation alignment has proven effective in prior works [yu2024representation, wu2025representation, jiang2025no, stracke2025cleandift], we specifically leverage the timestep-conditioned projection head adopted by CleanDIFT [stracke2025cleandift] and shown in Fig 3 (a). This projection head serves to project the single, noise-free representation $r_{D^{3}}^{K}$ into $T$ distinct timestep-dependent representation spaces for the purpose of matching the expert representations. This mechanism implicitly requires our model’s representation to encode the diverse diffusion priors corresponding to different noise levels. We measure the alignment by their point-wise distance:
@@ 3.2.2 Task-Specific Supervision
We meanwhile apply task-specific supervision by minimizing the task-specific loss $L_{\text{task}}$ to the D³-Predictor to leverage this aggregated diffusion prior for dense prediction tasks. For depth estimation, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss, the affine-invariant loss [ranftl2020towards], and the gradient loss in log space [ranftl2020towards]. For surface normal estimation, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss and the angular loss. For image matting, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss, $\ell_{1}$ loss the and gradient loss. Notably, the task-specific supervision and the self-supervised diffusion priors aggregation are jointly optimized. As a result, the aggregated diffusion prior preserves more task-relevant geometric structure information while suppressing irrelevant color-related information. Finally, the overall training objective $L$ is defined as:
@@ 3.3 Training on Mixed Data
The quality of real dense prediction datasets suffer from the physical limitations of sensing devices [huang2023neural]. Recent dense prediction methods [ke2024repurposing, zhao2025diception, xu2024matters] have instead opted to train entirely on synthetic datasets. However, we argue that the domain gap between synthetic and real data can hinder real-world generalization. In this work, we train our model on a balanced mixture of synthetic data and real images with pseudo labels. The synthetic data provide fine-grained geometric details, whereas the real images contribute to the model’s generalization by providing more diverse scenes.
Table 1: Quantitative comparison of different depth estimation methods across multiple zero-shot benchmarks. We use the same evaluation protocol as DICEPTION [zhao2025diception]. NFEs is the numbers of function evaluations (ensemble $\times$ steps)
@@
Table 2: Quantitative comparison of different surface normal estimation methods across multiple zero-shot benchmarks. We use the same evaluation protocol as Marigold-Normals [ke2025marigold]. NFEs is the numbers of function evaluations (ensemble $\times$ steps).
@@ 4.1.1 Training Datasets
We train D³-Predictor on different datasets for different dense prediction tasks: 1) For depth estimation: we use $10,777$ samples from Hypersim [roberts2021hypersim], a synthetic indoor scene dataset; $4,252$ samples from Virtual KITTI [cabon2020virtual], a synthetic driving scene dataset; and $15,160$ real images from COCO [lin2014microsoft] with pseudo-depth labels generated by Depth Pro [bochkovskii2024depth]. 2) For surface normal estimation: we use $29,284$ samples from Hypersim [roberts2021hypersim] and $376$ samples from Sintel [butler2012naturalistic], which contains both indoor and outdoor scenes from a short animated film. 3) For image matting: we use the training set of P3M-10K [li2021privacy], which includes $9,421$ images and is currently the largest portrait matting dataset. We only employ real images with pseudo labels for depth estimation, since other tasks have been less explored and their models generalize poorly to in-the-wild images. More details on the training data (e.g. data mixing strategies) are in Appendix.
@@ 4.1.2 Implementation Details
Unless otherwise specified, we initialize our D³-Predictor with Stable Diffusion v2.1 [rombach2022high] for fair comparison with previous works that also adopt it. We fine-tune the U-Net of it using an image resolution of $768\times 768$. We employ the Adam optimizer [loshchilov2017decoupled] with a batch size of 32 and a learning rate of 1e51e-5, following a linear warm-up schedule. All experiments are conducted on two NVIDIA L40S GPUs.

### v4

@@ Abstract
Although diffusion models with strong visual priors have emerged as powerful dense prediction backbones, they overlook a core limitation: the stochastic noise at the core of diffusion sampling is inherently misaligned with dense prediction that requires a deterministic mapping from image to geometry. In this paper, we show that this stochastic noise corrupts fine-grained spatial cues and pushes the model toward timestep-specific noise objectives, consequently destroying meaningful geometric structure mappings. To address this, we introduce D³-Predictor, a noise-free deterministic diffusion-based dense prediction model built by reformulating a pretrained diffusion model without stochasticity noise. Instead of relying on noisy inputs to leverage diffusion priors, D³-Predictor views the pretrained diffusion network as an ensemble of timestep-dependent visual experts and self-supervisedly aggregates their heterogeneous priors into a single, clean, and complete geometric prior. Meanwhile, we utilize task-specific supervision to seamlessly adapt this noise-free prior to dense prediction tasks. Extensive experiments on various dense prediction tasks demonstrate that D³-Predictor achieves competitive or state-of-the-art performance in diverse scenarios. In addition, it requires less than half the training data previously used and efficiently performs inference in a single step. Our code, data, and checkpoints are publicly available at https://x-gengroup.github.io/HomePage_D3-Predictor/.
@@ 1 Introduction
Dense prediction, such as depth estimation [50] and surface normal estimation [76], are fundamental tasks in computer vision, with numerous applications like autonomous driving [6, 20], scene reconstruction [10, 90], inverse rendering [9, 87], and so on. Although state-of-the-art discriminative dense prediction models [4, 25] achieve impressive performance, they still struggle to capture fine-grained high-frequency details. To address this limitation, current works [31, 41, 17, 84, 92] reformulate dense prediction as an image-conditioned iterative denoising process based on diffusion models [66]. By leveraging powerful visual priors of diffusion models, these methods can produce dense predictions results with fine-grained geometric details.
@@
While these diffusion-based dense prediction methods have demonstrated promising results, they still suffer from the stochastic noise inherent in diffusion models. Stochastic noise is a fundamental component of diffusion models, enabling sample diversity that is particularly beneficial for creative image [44, 79, 58] and video [83, 91, 75] generation. However, dense prediction is intrinsically deterministic, indicating that the stochastic noise essential to diffusion models may be misaligned with the task’s deterministic objective. We posit that the stochastic noise introduces two critical issues for dense prediction tasks: 1) Stochastic noise disrupts the geometric structures and small-scale objects in the input image, thereby degrading the integrity of input information and impeding precise spatial perception (cf. Fig. 2 (a)); 2) The stochastic noise drives diffusion models to focus on modeling noise distributions [24] instead of establishing the geometric structure mappings essential for dense prediction tasks [17]. Moreover, the iterative denoising process of the diffusion model further incurs substantial inference overhead. These observations motivate a fundamental question: Can diffusion models be reformulated into a noise-free deterministic framework to better suit dense prediction tasks?
@@
Recent works attempt to employ deterministic noise to alleviate the stochasticity in diffusion-based dense prediction methods. For example, GenPercept [80] and E2E-FT [18] eliminate stochasticity by fixing the noise schedule (a function of the timestep) to introduce deterministic noise. However, diffusion models learn timestep-specific objectives [2, 3], leading to distinct diffusion priors at each timestep. Consequently, fixing the noise schedule disrupts this prior structure, resulting in incomplete priors and a loss of geometric fidelity (cf. Fig. 2 (b)). On the other hand, StableNormal [84] suppresses stochasticity while preserving diffusion priors via a more complex two-stage pipeline with external DINO [55] guidance. However, this design incurs a higher computational overhead.
@@
In this work, we aim to fully eliminate the adverse effects of stochastic noise on diffusion-based dense prediction, without compromising the diffusion prior and with minimal additional computational cost. To this end, we propose D³-Predictor, a noise-free deterministic diffusion-based dense prediction model initialized from the pretrained diffusion model. Specifically, we treat the pretrained diffusion models at different timesteps as an ensemble of visual experts following the CleanDIFT [70] paradigm, each exhibiting distinct timestep-dependent diffusion priors. In this context, each visual expert takes a noisy image together with its timestep as input, while our D³-Predictor operates directly on a clean image. D³-Predictor then aggregates diffusion priors from multiple visual experts into a complete and noise-free one in a lightweight self-supervised manner, by aligning its internal representations with those of visual experts. We simultaneously apply task-specific supervision to the D³-Predictor to easily leverage this aggregated diffusion prior for various dense prediction tasks.
@@
  • We introduce D³-Predictor, an end-to-end noise-free deterministic diffusion-based dense prediction model. D³-Predictor combines self-supervised noise-free diffusion prior aggregation with task-specific supervision, allowing the model to transfer the diffusion priors into a noise-free geometric prior specifically for dense prediction tasks.
@@ 2.1 Diffusion-based Dense Prediction
Several works have explored leveraging diffusion models for dense prediction. Methods such as DIFT [71] and CleanDIFT [70] extract diffusion priors as general-purpose visual features and train lightweight probes for dense prediction tasks, but they are not end-to-end and their performance remains limited. While CleanDIFT [70] obtains noise-free general-purpose features, our method integrates this mechanism with task-specific supervision in an end-to-end manner. This joint optimization allows us to effectively transform the aggregated diffusion priors into noise-free geometric priors specifically tailored for dense prediction. On the other hand, other methods [31, 21, 92, 17] fine-tune a pretrained diffusion model for dense prediction tasks, achieving impressive results. However, these fine-tuning methods rely on stochastic noise to exploit diffusion priors. GenPercept [80] and E2E-FT [18] attempt to alleviate stochasticity but disrupt the integrity of diffusion priors, while StableNormal [84] suppresses stochasticity at the cost of additional computational overhead. In contrast, our method reformulates the diffusion model into a noise-free framework to better suit dense prediction in a simple self-supervised manner, ensuring both prediction performance and speed.
@@ 2.2 Self-Supervised Enhanced Diffusion Models
Recent studies [86, 78, 28] have shown that enhancing the internal representations of diffusion models, which serve as diffusion priors, can accelerate generative training and improve generation quality. Unlike methods [86] that rely on external pretrained features (e.g., DINO [55]), SRA [28] improves diffusion representations in a self-supervised way via self-representation alignment. The self-supervised enhancement in SRA [28] leads to improvements in both image generation quality and efficiency, demonstrating that diffusion models can provide representation guidance by themselves. In this work, we extend the idea of this self-supervised enhancement from diffusion-based generation to diffusion-based dense prediction. We align the noise-free internal representations of the diffusion model with their noisy counterparts at different noise levels in a self-supervised manner. This process produces complete, noise-free diffusion priors that are inherently more suitable for dense prediction.
@@ 3.1 Preliminary: Diffusion-based Dense Prediction
As illustrated in Fig. 3 (b), previous diffusion-based dense prediction methods, such as Marigold [31], adopt the time-consuming iterative denoising process described above, where the conditional input $c$ is an RGB image. At each iteration, the model progressively removes stochastic noise associated with a specific noise level and timestep, thus necessitating multi-step inference.
In contrast, recent work GenPercept [80] simplifies this process by fixing the noise schedule. However, it is worth noting that diffusion models have different objectives at different noise levels [2, 3]. At higher noise levels, the model captures coarse structural information, while at lower noise levels, it focuses on details [64]. Thus, the model learns distinct diffusion priors across timesteps, which together constitute a complete diffusion prior. Fixing the noise schedule disrupts this prior structure, leading to incomplete priors and suboptimal performance.
@@ 3.2.1 Self-Supervised Diffusion Priors Aggregation
The core of our approach is to train our D³-Predictor $f_{D^{3}}(\cdot)$ to incorporate a noise-free and complete diffusion prior compared to the pre-trained diffusion model, which is enhanced to be more suitable for deterministic dense prediction tasks. We achieve this enhancement through a simple self-supervised manner, and crucially, relying solely on the model itself rather than costly external guidance (e.g., DINO [55]) adapted by previous works [84]. We begin by initializing D³-Predictor with a pretrained diffusion model.
As shown in Fig 3 (a), the insight here is to view the diffusion model as an ensemble of $t$ visual experts, where each expert $f_{\text{exp}}(\cdot,\epsilon,t)$ exhibits a distinct timestep-dependent diffusion prior. We inherit the mechanism of aligning a clean backbone with a frozen diffusion model from CleanDIFT [70]. While the visual experts take a noisy image $(\epsilon,x)$ together with its timestep $t$ as input, our D³-Predictor operates directly on a clean image $x$. During training, D³-Predictor aligns its internal representations $r_{D^{3}}^{K}$ = $f_{D^{3}}^{K}(x)$ with those of $t$ visual experts $r_{\text{exp}}^{K,t}$ = $f_{\text{exp}}^{K}(x,\epsilon,t)$. Here, $K$ is the index set of model layers selected for alignment.
To stabilize this timestep-dependent representation alignment, we introduce a lightweight, trainable timestep-conditioned projection head $P_{\varphi}^{K}(\cdot,t)$ parameterized by $\varphi$. While the use of projection heads for representation alignment has proven effective in prior works [86, 78, 28, 70], we specifically leverage the timestep-conditioned projection head adopted by CleanDIFT [70] and shown in Fig 3 (a). This projection head serves to project the single, noise-free representation $r_{D^{3}}^{K}$ into $T$ distinct timestep-dependent representation spaces for the purpose of matching the expert representations. This mechanism implicitly requires our model’s representation to encode the diverse diffusion priors corresponding to different noise levels. We measure the alignment by their point-wise distance:
@@ 3.2.2 Task-Specific Supervision
We meanwhile apply task-specific supervision by minimizing the task-specific loss $L_{\text{task}}$ to the D³-Predictor to leverage this aggregated diffusion prior for dense prediction tasks. For depth estimation, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss, the affine-invariant loss [61], and the gradient loss in log space [61]. For surface normal estimation, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss and the angular loss. For image matting, $L_{\text{task}}$ is defined as a weighted sum of the MSE loss, $\ell_{1}$ loss the and gradient loss. Notably, the task-specific supervision and the self-supervised diffusion priors aggregation are jointly optimized. As a result, the aggregated diffusion prior preserves more task-relevant geometric structure information while suppressing irrelevant color-related information. Finally, the overall training objective $L$ is defined as:
@@ 3.3 Training on Mixed Data
The quality of real dense prediction datasets suffer from the physical limitations of sensing devices [26]. Recent dense prediction methods [31, 92, 80] have instead opted to train entirely on synthetic datasets. However, we argue that the domain gap between synthetic and real data can hinder real-world generalization. In this work, we train our model on a balanced mixture of synthetic data and real images with pseudo labels. The synthetic data provide fine-grained geometric details, whereas the real images contribute to the model’s generalization by providing more diverse scenes.
Table 1: Quantitative comparison of different depth estimation methods across multiple zero-shot benchmarks. We use the same evaluation protocol as DICEPTION [92]. NFEs is the numbers of function evaluations (ensemble $\times$ steps)
@@
Table 2: Quantitative comparison of different surface normal estimation methods across multiple zero-shot benchmarks. We use the same evaluation protocol as Marigold-Normals [32]. NFEs is the numbers of function evaluations (ensemble $\times$ steps).
@@ 4.1.1 Training Datasets
We train D³-Predictor on different datasets for different dense prediction tasks: 1) For depth estimation: we use $10,777$ samples from Hypersim [65], a synthetic indoor scene dataset; $4,252$ samples from Virtual KITTI [6], a synthetic driving scene dataset; and $15,160$ real images from COCO [45] with pseudo-depth labels generated by Depth Pro [4]. 2) For surface normal estimation: we use $29,284$ samples from Hypersim [65] and $376$ samples from Sintel [5], which contains both indoor and outdoor scenes from a short animated film. 3) For image matting: we use the training set of P3M-10K [43], which includes $9,421$ images and is currently the largest portrait matting dataset. We only employ real images with pseudo labels for depth estimation, since other tasks have been less explored and their models generalize poorly to in-the-wild images. More details on the training data (e.g. data mixing strategies) are in Appendix.
@@ 4.1.2 Implementation Details
Unless otherwise specified, we initialize our D³-Predictor with Stable Diffusion v2.1 [66] for fair comparison with previous works that also adopt it. We fine-tune the U-Net of it using an image resolution of $768\times 768$. We employ the Adam optimizer [46] with a batch size of 32 and a learning rate of 1e51e-5, following a linear warm-up schedule. All experiments are conducted on two NVIDIA L40S GPUs.